OpenStudy (moongazer):

Find the speed of the car in miles per hour, if the car makes 120 rev/min. and the radius 24 in.

OpenStudy (moongazer):

I got 17.136 miles per hour am I correct?

OpenStudy (lgbasallote):

i got the same...

OpenStudy (lgbasallote):

\[120 \times 2 \times \pi \times 24 = 18,095 in/min \implies 17. 14 mph\]

OpenStudy (moongazer):

thanks :) I used linear speed and angular speed:)

OpenStudy (lgbasallote):

i used google converter....lol jk

OpenStudy (moongazer):

lol